OpenStudy (anonymous):

What is 18 ÷ 2/3? A. 1/12 B. 3 C. 12 D. 27

OpenStudy (anonymous):

is it d?

OpenStudy (anonymous):

18 ÷ 2/3=18 * 3/2 = 27 and yea !

OpenStudy (aakashsudhakar):

Dividing by a fraction is the same as multiplying by its reciprocal. Therefore, 18/(2/3) is the same as 18 * (3/2). 18 times 3 gets you 54/2, which simplifies to 27. So yes, the answer is (D)!

OpenStudy (anonymous):

You are right @studyislandhelp it is D

OpenStudy (anonymous):

i knew it!

OpenStudy (anonymous):

answer: D